Find each of the function below, indicate whether the function in onto, on-to-one neither or both. If the function is not onto or nor one-to-one, give an example showing why

G;R      R. g(x)=x^3


1
Expert's answer
2021-07-06T08:54:36-0400

Consider the function "g:\R\to\R,\ g(x)=x^3." Let "g(x_1)=g(x_2)." Then "x_1^3=x_2^3," and hence "x_1=x_2." We conclude that this function is one-to-one. For any "y\in\R" the equation "x^3=y" has a unique solution "x=\sqrt[3]{y}," and hence for any "y\in\R" there exists "x\in \R" such that "f(x)=y." It follows that the function "f" is onto. Therefore, this function is a bijection.